1. Determine the power requirement to pull a four bottom 30 cm mould board plough working to a depth of 15 cm.


The tractor is operated at a speed of 6 km/h. The soil resistance is 0.7 kg/cm2


1
Expert's answer
2021-06-21T11:42:19-0400

"h=4h_0=4\cdot 30=120~cm,"

"A=hl=120\cdot 15=1800~cm^2,"

"F=pAg=0.7\cdot 1800\cdot 9.8=12348~N,"

"P=Fv=12348\cdot \frac{6}{3.6}=20.58~kW."